How many words of length 10 can you form where no letter is repeated?
1
Expert's answer
2020-05-20T18:32:43-0400

Solution.

Since the Italian alphabet contains 21 letters, the number of words of 10 letters can be found:

"A_n^k=\dfrac{n!}{(n-m)!} -" Placement from n to k ;

"A_{21}^{10}=\\dfrac{21!}{(21-10)!}=\\dfrac{21!}{11!}=12\\sdot13\\sdot14\\sdot15\\sdot16\\sdot17\\sdot18\\sdot19\\sdot\\ 20 \\sdot21=\n\u202c"

"=1\u00a0279\u00a0935\u00a0820\u00a0800\u202c;"

Answer: "A_{21}^{10}=1\u00a0279\u00a0935\u00a0820\u00a0800\u202c."